# ISRE-0001-2024.00

### v1

@@ 2.2 Label Bias
Label bias has gained substantial attention as a key issue affecting data quality. Label bias has been defined as a systematic discrepancy between the outcome of interest and the observed labels, such that the relationship underlying the mismatch varies across groups  (Li et al. 2022, 2024). Different sources of label bias have been identified (Li et al. 2022, 2024). One significant source of label bias is the human labeling bias, which arises when labels reflect annotators’ cognitive biases, prejudices, or limited domain knowledge (Howe 2008). With the rise of crowd-sourcing services (Howe 2008), researchers have increasingly highlighted the risks posed by annotator cognitive biases (Eickhoff 2018, Draws et al. 2021) and the potential presence of prejudice in labeling decisions (Otterbacher 2015). For example, hate speech annotators’ judgments are influenced by biased assumptions in language, leading to unfair and disproportionate impacts on marginalized groups  (Davani et al. 2023). Expert assessments can also embed biases. In healthcare, for example, it has been shown that the quality of pain assessment and treatment recommendations can be undermined by provider biases  (Hoffman et al. 2016). Another major source of label bias is label measurement bias, which arises when the label being measured aligns conceptually with the construct of interest but is systematically mismeasured across different groups  (Jacobs and Wallach 2021). This type of bias is prevalent when measurement tools are designed around the needs of one group, overlooking others. A clear example is pulse oximetry, which has been shown to overestimate oxygen saturation levels in patients with darker skin tones. Black patients experience occult hypoxemia—low oxygen levels undetected by pulse oximetry—at nearly three times the rate of white patients (Sjoding et al. 2020).
@@
Label bias can also arise from a label construct gap, which occurs when there is a discrepancy in theoretical definitions between the construct of interest and the construct observed (Jacobs and Wallach 2021, Passi and Barocas 2019, Li et al. 2022). This discrepancy often stems from the repurposing of pre-existing datasets, where the original data collection goals differ from the intended task (Agarwal and Dhar 2014). For example, in healthcare, insurance claim costs often serve as a proxy for patient needs, but due to historical disparities Black patients incur lower expenses despite equal or greater medical needs, leading to inaccurate assessments  (Obermeyer et al. 2019). Similarly, in college admissions, GPA or class rankings may be used as proxies for student success, ignoring broader factors such as leadership, creativity, and resilience, thereby disadvantaging certain demographics  (Suresh and Guttag 2021).
@@
Despite its widespread presence, research focused on label bias has primarily centered on characterizing or conceptualizing it. Li et al. (2022) empirically demonstrates that collecting more data can exacerbate bias if label bias is overlooked. Fogliato et al. (2020) finds that even small biases in observed labels can produce disparate performance across races of recidivism risk assessment tools, and Akpinar et al. (2021) shows that differential rates in crime reporting can lead to bias in predictive policing systems. DeCoLe advances this literature by moving beyond characterizing label bias to detecting mislabeled instances in datasets potentially affected by it, providing an algorithmic solution to perform bias-aware mislabeling detection that supports effective mitigation of label bias at scale.
@@ 2.3 Bias Mitigation in Artificial Intelligence (AI)
The risk that bias embedded in data may yield biased AI systems have gained significant attention in the past years (Barocas and Selbst 2016, De-Arteaga et al. 2022). This has given rise to a set of methodologies to mitigate bias of different forms (De-Arteaga et al. 2022). Existing approaches have sought to address the underrepresentation of certain groups through approaches such as adaptive instance selection (Cai et al. 2022), synthetic data augmentation (Theodorou et al. 2025), or statistical sampling correction (Zelaya et al. 2019). Another body of work has sought to modify the feature space in order to remove unwanted correlations with protected attributes (Zemel et al. 2013, Caliskan et al. 2017, Bolukbasi et al. 2016, Zhao et al. 2018, Bolukbasi et al. 2016, Yu et al. 2023, Blodgett et al. 2020). In the context of natural language processing, research has demonstrated that word embedding — vector representations of words capturing their meanings and relationships— encodes associations that reflect and reinforce societal biases (Caliskan et al. 2017, Bolukbasi et al. 2016). To address this, researchers have developed strategies that aim to de-bias embeddings while preserving linguistic integrity (Zhao et al. 2018, Bolukbasi et al. 2016, Yu et al. 2023, Blodgett et al. 2020). While these efforts address representation issues and feature space bias in data, our work complements this line of research by focusing on label bias.
@@ 2.3.1 Learning from Erroneous and Biased Labels
A stream of literature on supervised machine learning has aimed to develop methodologies that are robust to errors in the labels used to train them. This includes methods that aggregate labels from multiple annotators using probabilistic models  (Zhang et al. 2016, Snow et al. 2008, Smyth et al. 1994, Dawid and Skene 1979, Whitehill et al. 2009, Welinder et al. 2010, Yan et al. 2010, Zhang et al. 2016), approaches that train models directly on noisy labels by adjusting loss functions or model architectures (Shu et al. 2019, Natarajan et al. 2013, Ma et al. 2020, Lipton et al. 2018, Van Rooyen et al. 2015, Patrini et al. 2017, Chen et al. 2019), and techniques that manage different misclassification costs via post hoc decision threshold adjustments (Elkan 2001).
@@
The risk of label bias—and in particular the possible presence of group- and class-conditional label errors—has been recognized as a problem by this stream of research, and a few works have proposed methodologies to learn predictive models when such bias is present in the data (Wang et al. 2021, Jiang and Nachum 2020, Kamiran and Calders 2009). These methods have introduced techniques that modify loss functions to enable robust learning in the presence of these types of errors. Wang et al. (2021) propose fairness-aware surrogate loss functions and constraints. However, the method relies on knowing the group- and class-conditional error rates a priori, which is information that is typically inaccessible in practice. Jiang and Nachum (2020) develop an iterative re-weighting method to satisfy fairness constraints, but the selection of instances for up-weighting or down-weighting is based solely on group membership and observed labels, without identifying which instances are actually mislabeled. As a result, the method cannot be used to improve the quality of the data itself. Kamiran and Calders (2009) propose identifying the minimum data perturbation required to enforce or approximate demographic parity, a fairness criterion that assumes equal label distributions across groups. However, this assumption is often violated in real-world contexts where outcome rates differ naturally across populations. In healthcare, for example, certain diseases are more prevalent among specific demographic groups; in social services, some populations are at higher risk of adverse outcomes; and in online content moderation, hate speech may disproportionately target some groups. Crucially, all these methods pursue specific fairness objectives evaluated at the inference stage and focus on mitigating the consequences of label bias for supervised machine learning tasks. In contrast, DeCoLe addresses label bias directly by performing bias-aware mislabeling detection within the data itself, making it a critical component in facilitating label bias mitigation and data quality improvement. This upstream, data-centric approach not only complements model-level fairness interventions but also improves data quality in a way that supports a broader range of tasks beyond supervised learning.
@@ 3.1 Bias-Aware Mislabeling Detection
We consider the problem of detecting mislabeled instances in a dataset that may be affected by label bias, a task we refer to as bias-aware mislabeling detection. Let $\bm{D}\coloneqq(\bm{x},g,\tilde{y})^{n}$ denote a dataset of $n$ instances, where each instances is represented by features $\bm{x}$ with associated observed label $\tilde{y}\in\{0,1\}$ and group membership $g$. Group membership may denote attributes such as age, gender, race, or an intersection of multiple of these. Finally, let $y^{*}$ represent a latent gold standard label.
We assume there exists a group-and class-conditional noise such that there is a non-correspondence between the observed label $\tilde{y}$ and the latent gold standard label $y^{*}$ that depends on both the gold standard label, $y^{*}$, and the group membership, $g$. Let $\pi_{c,g_{k}}$ denote the misclassification rate for observed labels $\tilde{y}=c$. Specifically, for a particular group $g_{k}$, where $k\in\{0,...,(K-1)\}$ and $K$ denotes the total number of possible groups, positive instances may be misclassified as negative with a probability $\pi_{0,{g_{k}}}=p(\tilde{y}=0|y^{*}=1,g=g_{k})$, and negative instances may be misclassified as positive with a probability $\pi_{1,{g_{k}}}=p(\tilde{y}=1|y^{*}=0,g=g_{k})$. Label bias occurs when there is a disparity in one (or both) of these metrics across different groups $g_{k}$  (Li et al. 2022). We formulate the problem under a binary labeling setting, while allowing for multi-categorical group memberships. Our goal is to provide a generic framework that can be used to identify mislabeled instances in a dataset that is potentially affected by label bias.
@@
$\bm{D}_{\tilde{y}=i,y^{*}=j}$ The set of instances such that $\tilde{y}=i$ and $y^{*}=j$.
$\bm{D}_{g_{k}}$ The set of instances belonging to group $g_{k}$.
@@
$\bm{X}_{g_{k}}$ The set of feature vectors for all instances in group $g_{k}$.
$\hat{p}(\bm{x}_{g_{k}})$ The probability predicted by classifier $f_{k}$ for $\bm{x}_{g_{k}}\in\bm{X}_{g_{k}}$ to be in the positive class.
@@
$\bm{D}_{\tilde{y}\neq y^{*}|g}$ Mislabeled instances.
$\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ Instances estimated to be mislabeled by DeCoLe applied considering group $g$.
$\bm{D}_{\tilde{y}=y^{*}|g}$ Correctly labeled instances.
$\bm{\hat{D}}_{\tilde{y}=y^{*}|g}$ Instances estimated to be correctly labeled by DeCoLe applied considering group $g$.
@@ 3.2 Decoupled Confident Learning (DeCoLe)
We propose Decoupled Confident Learning (DeCoLe) to effectively perform bias-aware mislabeling detection. DeCoLe only needs access to the observed dataset $\bm{D}\coloneqq(\bm{x},g,\tilde{y})^{n}$, and does not require access to any gold standard labels $y^{*}$. The input to DeCoLe is the original data, $\bm{D}$, and the output is a partition of the data into two sets, one estimated to be mislabeled, and the other one estimated to be correctly labeled, as shown in Figure 1.
@@
At a high level, DeCoLe addresses the challenge of bias-aware mislabeling detection by separately performing a series of confident learning procedures for each group. This group-specific approach enables independent estimation of the joint distribution between observed biased labels and latent gold standard labels, effectively detecting label errors under group- and class-conditional noise. In doing this, DeCoLe builds on the common machine learning principle that predicted probabilities reflect confidence in classifications. This idea is formalized in the concept of “probably approximately correct identification”  (Valiant 1984), which, within the theory of learnability, states that after randomly observing both positive and negative instances, an identification procedure should infer a hypothesis that, with high probability, is close to the correct concept (Angluin and Laird 1988, Valiant 1984). In our approach, we incorporate the insight that the underlying concept may differ across groups  (Dwork et al. 2018, Gupta et al. 2023). As a result, effective identification requires inferring group-specific hypotheses, each guided by its own tailored identification procedure.
@@
    Train a separate predictive model $f_{k}$ for each group $g_{k}$, and obtain out-of-sample predicted probabilities $\hat{p}(\bm{x}_{g_{k}})=\hat{p}(\tilde{y}=1;\bm{x}_{g_{k}},f_{k})\>\forall\>x_{g_{k}}\in\bm{D}$ , where $\bm{x}_{g_{k}}$ denotes $\bm{x}$ for $(\bm{x},\tilde{y})\in\bm{D}_{g_{k}}$.This process yields predicted probabilities $\hat{p}(\bm{x})$ for every instance in the dataset.
@@
    $\text{LB}_{g_{k}}=\frac{1}{|D_{\tilde{y}=1,g=g_{k}}|}\sum_{\bm{x}_{g_{k}}\in\bm{D}_{\tilde{y}=1}}\hat{p}(\bm{x}_{g_{k}}),$ (1)
    $\text{UB}_{g_{k}}=\frac{1}{|D_{\tilde{y}=0,g=g_{k}}|}\sum_{\bm{x}_{g_{k}}\in\bm{D}_{\tilde{y}=0}}\hat{p}(\bm{x}_{g_{k}}).$ (2)
@@
    $\text{CPS}_{g_{k}}=\left\{(\bm{x},\tilde{y})\in\bm{D}_{g_{k}}:\hat{p}(\bm{x})\geq\text{LB}_{g_{k}}\right\}$ (3)
    $\text{CNS}_{g_{k}}=\left\{(\bm{x},\tilde{y})\in\bm{D}_{g_{k}}:\hat{p}(\bm{x})\leq\text{UB}_{g_{k}}\right\}.$ (4)
@@
Input: A dataset with noisy labels $\bm{D}\coloneqq(\bm{x},g,\tilde{y})^{n}$, initialize a set of classifier $\{f_{1},...,f_{k}\}$.
@@
     Initialize a group-specific bias detection set $\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g=g_{k}}=\emptyset$
     Part 1: Estimating $p(\bm{x})$.
     $f_{k}$.fit$(\bm{D}_{g_{k}})$
     $\hat{p}(\bm{X}_{g_{k}})\leftarrow$ $f_{k}$.predict_crossval_prob $(\tilde{y}=1|\bm{X}_{g_{k}})$ where $\bm{X}_{g_{k}}:=\left\{\bm{x}:(\bm{x},\tilde{y})\in\bm{D}_{g_{k}}\right\}$
@@
$\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}\coloneqq\{\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g=g_{k}}\}^{K}$
return $\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}$
The detailed algorithm can be found in Algorithm 1. When there is group and class-conditional noise, decoupling—training individual models for each group—disentangles the group-specific noise structure, and accounts for differential subgroup validity (Gupta et al. 2023). By training separate classifiers for each group, DeCoLe allows the likelihood of an instance belonging to a specific class to be estimated in a way that reflects the unique characteristics and optimization needs of the group, reducing the tension between optimizing for one group at the expense of another. Importantly, such an approach allows us to accommodate different error patterns across groups: for example, one group may be more susceptible to incorrect observed labels in the form of false positives, while another group is more susceptible to incorrect observed labels in the form of false negatives.
@@
In the next section, we examine sufficient conditions where DeCoLe perfectly detects all mislabeled instances, even when $\hat{p}(\bm{x})$ is noisy.

### v2

@@ 2.2   Label Bias
Label bias has gained substantial attention as a key issue affecting data quality. Label bias has been defined as a systematic discrepancy between the outcome of interest and the observed labels, such that the relationship underlying the mismatch varies across groups  (li2022more, ; li2024label, ). Different sources of label bias have been identified (li2022more, ; li2024label, ). One significant source of label bias is the human labeling bias, which arises when labels reflect annotators’ cognitive biases, prejudices, or limited domain knowledge (howe2008crowdsourcing, ). With the rise of crowd-sourcing services (howe2008crowdsourcing, ), researchers have increasingly highlighted the risks posed by annotator cognitive biases (eickhoff2018cognitive, ; draws2021checklist, ) and the potential presence of prejudice in labeling decisions (otterbacher2015crowdsourcing, ). For example, hate speech annotators’ judgments are influenced by biased assumptions in language, leading to unfair and disproportionate impacts on marginalized groups  (davani2023hate, ). Expert assessments can also embed biases. In healthcare, for example, it has been shown that the quality of pain assessment and treatment recommendations can be undermined by provider biases  (hoffman2016racial, ). Another major source of label bias is label measurement bias, which arises when the label being measured aligns conceptually with the construct of interest but is systematically mismeasured across different groups  (jacobs2021measurement, ). This type of bias is prevalent when measurement tools are designed around the needs of one group, overlooking others. A clear example is pulse oximetry, which has been shown to overestimate oxygen saturation levels in patients with darker skin tones. Black patients experience occult hypoxemia—low oxygen levels undetected by pulse oximetry—at nearly three times the rate of white patients (sjoding2020racial, ).
@@
Label bias can also arise from a label construct gap, which occurs when there is a discrepancy in theoretical definitions between the construct of interest and the construct observed (jacobs2021measurement, ; passi2019problem, ; li2022more, ). This discrepancy often stems from the repurposing of pre-existing datasets, where the original data collection goals differ from the intended task (agarwal2014big, ). For example, in healthcare, insurance claim costs often serve as a proxy for patient needs, but due to historical disparities Black patients incur lower expenses despite equal or greater medical needs, leading to inaccurate assessments  (obermeyer2019dissecting, ). Similarly, in college admissions, GPA or class rankings may be used as proxies for student success, ignoring broader factors such as leadership, creativity, and resilience, thereby disadvantaging certain demographics  (suresh2021framework, ).
@@
Despite its widespread presence, research focused on label bias has primarily centered on characterizing or conceptualizing it. li2022more empirically demonstrates that collecting more data can exacerbate bias if label bias is overlooked. fogliato2020fairness finds that even small biases in observed labels can produce disparate performance across races of recidivism risk assessment tools, and akpinar2021effect shows that differential rates in crime reporting can lead to bias in predictive policing systems. DeCoLe advances this literature by moving beyond characterizing label bias to detecting mislabeled instances in datasets potentially affected by it, providing an algorithmic solution to perform bias-aware mislabeling detection that supports effective mitigation of label bias at scale.
@@ 2.3   Bias Mitigation in Artificial Intelligence (AI)
The risk that bias embedded in data may yield biased AI systems have gained significant attention in the past years (barocas2016big, ; de2022algorithmic, ). This has given rise to a set of methodologies to mitigate bias of different forms (de2022algorithmic, ). Existing approaches have sought to address the underrepresentation of certain groups through approaches such as adaptive instance selection (cai2022adaptive, ), synthetic data augmentation (theodorou2025improving, ), or statistical sampling correction (zelaya2019parametrised, ). Another body of work has sought to modify the feature space in order to remove unwanted correlations with protected attributes (zemel2013learning, ; caliskan2017semantics, ; bolukbasi2016man, ; zhao2018learning, ; bolukbasi2016man, ; yu2023unlearning, ; blodgett2020language, ). In the context of natural language processing, research has demonstrated that word embedding — vector representations of words capturing their meanings and relationships— encodes associations that reflect and reinforce societal biases (caliskan2017semantics, ; bolukbasi2016man, ). To address this, researchers have developed strategies that aim to de-bias embeddings while preserving linguistic integrity (zhao2018learning, ; bolukbasi2016man, ; yu2023unlearning, ; blodgett2020language, ). While these efforts address representation issues and feature space bias in data, our work complements this line of research by focusing on label bias.
@@ 2.3.1   Learning from Erroneous and Biased Labels
A stream of literature on supervised machine learning has aimed to develop methodologies that are robust to errors in the labels used to train them. This includes methods that aggregate labels from multiple annotators using probabilistic models  (zhang2016learning, ; snow2008cheap, ; smyth1994inferring, ; dawid1979maximum, ; whitehill2009whose, ; welinder2010multidimensional, ; yan2010modeling, ; zhang2016learning, ), approaches that train models directly on noisy labels by adjusting loss functions or model architectures (shu2019meta, ; natarajan2013learning, ; ma2020normalized, ; lipton2018detecting, ; van2015learning, ; patrini2017making, ; chen2019understanding, ), and techniques that manage different misclassification costs via post hoc decision threshold adjustments (elkan2001foundations, ).
@@
The risk of label bias—and in particular the possible presence of group- and class-conditional label errors—has been recognized as a problem by this stream of research, and a few works have proposed methodologies to learn predictive models when such bias is present in the data (wang2021fair, ; jiang2020identifying, ; kamiran2009classifying, ). These methods have introduced techniques that modify loss functions to enable robust learning in the presence of these types of errors. wang2021fair propose fairness-aware surrogate loss functions and constraints. However, the method relies on knowing the group- and class-conditional error rates a priori, which is information that is typically inaccessible in practice. jiang2020identifying develop an iterative re-weighting method to satisfy fairness constraints, but the selection of instances for up-weighting or down-weighting is based solely on group membership and observed labels, without identifying which instances are actually mislabeled. As a result, the method cannot be used to improve the quality of the data itself. kamiran2009classifying propose identifying the minimum data perturbation required to enforce or approximate demographic parity, a fairness criterion that assumes equal label distributions across groups. However, this assumption is often violated in real-world contexts where outcome rates differ naturally across populations. In healthcare, for example, certain diseases are more prevalent among specific demographic groups; in social services, some populations are at higher risk of adverse outcomes; and in online content moderation, hate speech may disproportionately target some groups. Crucially, all these methods pursue specific fairness objectives evaluated at the inference stage and focus on mitigating the consequences of label bias for supervised machine learning tasks. In contrast, DeCoLe addresses label bias directly by performing bias-aware mislabeling detection within the data itself, making it a critical component in facilitating label bias mitigation and data quality improvement. This upstream, data-centric approach not only complements model-level fairness interventions but also improves data quality in a way that supports a broader range of tasks beyond supervised learning.
@@ 3.1   Bias-Aware Mislabeling Detection
We consider the problem of detecting mislabeled instances in a dataset that may be affected by label bias, a task we refer to as bias-aware mislabeling detection. Let $\boldsymbol{D}\coloneqq(\boldsymbol{x},g,\tilde{y})^{n}$ denote a dataset of $n$ instances, where each instances is represented by features $\boldsymbol{x}$ with associated observed label $\tilde{y}\in\{0,1\}$ and group membership $g$. Group membership may denote attributes such as age, gender, race, or an intersection of multiple of these. Finally, let $y^{*}$ represent a latent gold standard label.
We assume there exists a group-and class-conditional noise such that there is a non-correspondence between the observed label $\tilde{y}$ and the latent gold standard label $y^{*}$ that depends on both the gold standard label, $y^{*}$, and the group membership, $g$. Let $\pi_{c,g_{k}}$ denote the misclassification rate for observed labels $\tilde{y}=c$. Specifically, for a particular group $g_{k}$, where $k\in\{0,...,(K-1)\}$ and $K$ denotes the total number of possible groups, positive instances may be misclassified as negative with a probability $\pi_{0,{g_{k}}}=p(\tilde{y}=0|y^{*}=1,g=g_{k})$, and negative instances may be misclassified as positive with a probability $\pi_{1,{g_{k}}}=p(\tilde{y}=1|y^{*}=0,g=g_{k})$. Label bias occurs when there is a disparity in one (or both) of these metrics across different groups $g_{k}$  (li2022more, ). We formulate the problem under a binary labeling setting, while allowing for multi-categorical group memberships. Our goal is to provide a generic framework that can be used to identify mislabeled instances in a dataset that is potentially affected by label bias.
@@
$\boldsymbol{D}_{\tilde{y}=i,y^{*}=j}$ The set of instances such that $\tilde{y}=i$ and $y^{*}=j$.
$\boldsymbol{D}_{g_{k}}$ The set of instances belonging to group $g_{k}$.
@@
$\boldsymbol{X}_{g_{k}}$ The set of feature vectors for all instances in group $g_{k}$.
$\hat{p}(\boldsymbol{x}_{g_{k}})$ The probability predicted by classifier $f_{k}$ for $\boldsymbol{x}_{g_{k}}\in\boldsymbol{X}_{g_{k}}$ to be in the positive class.
@@
$\boldsymbol{D}_{\tilde{y}\neq y^{*}|g}$ Mislabeled instances.
$\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ Instances estimated to be mislabeled by DeCoLe applied considering group $g$.
$\boldsymbol{D}_{\tilde{y}=y^{*}|g}$ Correctly labeled instances.
$\boldsymbol{\hat{D}}_{\tilde{y}=y^{*}|g}$ Instances estimated to be correctly labeled by DeCoLe applied considering group $g$.
@@ 3.2   Decoupled Confident Learning (DeCoLe)
We propose Decoupled Confident Learning (DeCoLe) to effectively perform bias-aware mislabeling detection. DeCoLe only needs access to the observed dataset $\boldsymbol{D}\coloneqq(\boldsymbol{x},g,\tilde{y})^{n}$, and does not require access to any gold standard labels $y^{*}$. The input to DeCoLe is the original data, $\boldsymbol{D}$, and the output is a partition of the data into two sets, one estimated to be mislabeled, and the other one estimated to be correctly labeled, as shown in Figure 1.
@@
At a high level, DeCoLe addresses the challenge of bias-aware mislabeling detection by separately performing a series of confident learning procedures for each group. This group-specific approach enables independent estimation of the joint distribution between observed biased labels and latent gold standard labels, effectively detecting label errors under group- and class-conditional noise. In doing this, DeCoLe builds on the common machine learning principle that predicted probabilities reflect confidence in classifications. This idea is formalized in the concept of “probably approximately correct identification”  (valiant1984theory, ), which, within the theory of learnability, states that after randomly observing both positive and negative instances, an identification procedure should infer a hypothesis that, with high probability, is close to the correct concept (angluin1988learning, ; valiant1984theory, ). In our approach, we incorporate the insight that the underlying concept may differ across groups  (dwork2018decoupled, ; gupta2023same, ). As a result, effective identification requires inferring group-specific hypotheses, each guided by its own tailored identification procedure.
@@
    Train a separate predictive model $f_{k}$ for each group $g_{k}$, and obtain out-of-sample predicted probabilities $\hat{p}(\boldsymbol{x}_{g_{k}})=\hat{p}(\tilde{y}=1;\boldsymbol{x}_{g_{k}},f_{k})\>\forall\>x_{g_{k}}\in\boldsymbol{D}$ , where $\boldsymbol{x}_{g_{k}}$ denotes $\boldsymbol{x}$ for $(\boldsymbol{x},\tilde{y})\in\boldsymbol{D}_{g_{k}}$.This process yields predicted probabilities $\hat{p}(\boldsymbol{x})$ for every instance in the dataset.
@@
    $\text{LB}_{g_{k}}=\frac{1}{|D_{\tilde{y}=1,g=g_{k}}|}\sum_{\boldsymbol{x}_{g_{k}}\in\boldsymbol{D}_{\tilde{y}=1}}\hat{p}(\boldsymbol{x}_{g_{k}}),$ (1)
    $\text{UB}_{g_{k}}=\frac{1}{|D_{\tilde{y}=0,g=g_{k}}|}\sum_{\boldsymbol{x}_{g_{k}}\in\boldsymbol{D}_{\tilde{y}=0}}\hat{p}(\boldsymbol{x}_{g_{k}}).$ (2)
@@
    $\text{CPS}_{g_{k}}=\left\{(\boldsymbol{x},\tilde{y})\in\boldsymbol{D}_{g_{k}}:\hat{p}(\boldsymbol{x})\geq\text{LB}_{g_{k}}\right\}$ (3)
    $\text{CNS}_{g_{k}}=\left\{(\boldsymbol{x},\tilde{y})\in\boldsymbol{D}_{g_{k}}:\hat{p}(\boldsymbol{x})\leq\text{UB}_{g_{k}}\right\}.$ (4)
@@
Input: A dataset with noisy labels $\boldsymbol{D}\coloneqq(\boldsymbol{x},g,\tilde{y})^{n}$, initialize a set of classifier $\{f_{1},...,f_{k}\}$.
@@
     Initialize a group-specific bias detection set $\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g=g_{k}}=\emptyset$
     Part 1: Estimating $p(\boldsymbol{x})$.
     $f_{k}$.fit$(\boldsymbol{D}_{g_{k}})$
     $\hat{p}(\boldsymbol{X}_{g_{k}})\leftarrow$ $f_{k}$.predict_crossval_prob $(\tilde{y}=1|\boldsymbol{X}_{g_{k}})$ where $\boldsymbol{X}_{g_{k}}:=\left\{\boldsymbol{x}:(\boldsymbol{x},\tilde{y})\in\boldsymbol{D}_{g_{k}}\right\}$
@@
$\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}\coloneqq\{\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g=g_{k}}\}^{K}$
return $\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}$
The detailed algorithm can be found in Algorithm 1. When there is group and class-conditional noise, decoupling—training individual models for each group—disentangles the group-specific noise structure, and accounts for differential subgroup validity (gupta2023same, ). By training separate classifiers for each group, DeCoLe allows the likelihood of an instance belonging to a specific class to be estimated in a way that reflects the unique characteristics and optimization needs of the group, reducing the tension between optimizing for one group at the expense of another. Importantly, such an approach allows us to accommodate different error patterns across groups: for example, one group may be more susceptible to incorrect observed labels in the form of false positives, while another group is more susceptible to incorrect observed labels in the form of false negatives.
@@
In the next section, we examine sufficient conditions where DeCoLe perfectly detects all mislabeled instances, even when $\hat{p}(\boldsymbol{x})$ is noisy.